### 13 Mar 2024 v3

Abstract
The ongoing debate surrounding the impact of the Internet Research Agency’s (IRA) social media campaign during the 2016 U.S. presidential election has largely overshadowed the involvement of other actors. Our analysis brings to light a substantial group of suspended Twitter users, outnumbering the IRA user group by a factor of 60, who align with the ideologies of the IRA campaign. Our study demonstrates that this group of suspended Twitter accounts significantly influenced individuals categorized as undecided or weak supporters, potentially with the aim of swaying their opinions, as indicated by Granger causality.
keywords:
Social network, Disinformation, Election, Russian trolls
1 Introduction
Social media platforms have become increasingly prominent in shaping political events and social discussions. Political campaigns across the globe are heavily reliant on social media platforms to communicate with the masses and shape public opinion [1, 2, 3, 4, 5]. However, the rise of social media has also resulted in debates about their impact on society and the potential risks associated with their use.
Social media platforms, while holding the potential to facilitate communication and foster informed discussions, are also susceptible to the dissemination of misinformation and disinformation campaigns [6, 7, 8]. This issue extends beyond politics and seeps into sensitive domains like public health, as exemplified by the anti-vaccine movements during the COVID-19 pandemic [9]. Compelling evidence abounds, pointing to the active exploitation of social media platforms by certain governments to subvert domestic social movements and interfere in the democratic elections of foreign adversaries [10]. Noteworthy instances of such foreign interventions include the case of the 2017 French presidential election [11] and the highly significant interference by the Internet Research Agency (IRA: a Russian company engaged in online influence operations on behalf of Russian business and political interests) in the 2016 US presidential election [12, 13].
As outlined in the U.S. Special Counsel’s report [14], the Internet Research Agency initiated Russian interference operations as early as 2009. Their strategic approach involved the creation of social network campaigns aimed at fueling and magnifying political and social divisions within the United States [14, 15]. At the beginning of 2018, Twitter committed to the United States Congress and the public to provide regular updates and information regarding their investigation into foreign interference in U.S. political conversations on Twitter. In October 2018, Twitter openly released all the accounts and related content associated with potential information operations they had found on Twitter since 2016. This dataset consists of more than three thousand accounts affiliated with the IRA. It contains more than 9 million tweets, including the earliest Twitter activity of the accounts connected with these campaigns, dating back to 2009. The Twitter corporation estimates that 9% of the tweets from IRA accounts were election-related.
Since then, the number of works focusing on the role the IRA agency played in the 2016 US political campaign and social debates increased. A. Badawy et al.[16] found that conservatives retweeted Russian trolls significantly more often than liberals and produced 36 times more tweets. Among the 5.7 million distinct users analyzed between September 16 and November 9, 2016, about 4.9% and 6.2% of liberal and conservative users, respectively, were automated accounts (bots) used to share troll content. Text analysis of the content shared by trolls reveals that they had a mostly conservative, pro-Trump agenda. P. N. Howard et al. [17] concluded that the Russian strategies targeted many communities within the United States, particularly the most extreme conservatives and those with particular sensitivities to race and immigration. They found that IRA used a variety of fake accounts to infiltrate political discussions in liberal and conservative communities, including black activist communities, to exacerbate social divisions and influence the agenda. By combining network science and volumetric analysis, L. G. Stewart et al. found that troll accounts shared content to polarized information networks, likely accentuating disagreement and fostering division [18]. The conclusions above align with the findings of R. DiResta et al. [19], who observed that the IRA campaign was designed to exploit societal fractures, blur the lines between reality and fiction, and erode trust in media entities and the information environment, in government, in each other, and in democracy itself. In their study on disinformation S. Zannettou et al. [20], conducted an investigation into the behavioral differences between IRA and random Twitter users. The findings revealed that IRA users exhibit a higher tendency to disseminate content related to politics. Additionally, IRA employed multiple identities throughout the lifespan of their accounts and made deliberate efforts to amplify their impact on Twitter by increasing their number of followers.
The studies mentioned above aim to characterize the IRA campaign. In an attempt to evaluate the impact of the IRA campaign on Twitter users, C. Bail et al. conducted a study using a longitudinal survey that describes the attitudes and online behaviors of one thousand Republican and Democratic Twitter users in late 2017 [21]. Their findings suggest that Russian trolls might have failed to sow discord because they mostly interacted with those who were already highly polarized. In [22], N. Grinberg et al. demonstrated that exposure to fake news content during the 2016 elections was typically concentrated among a small group of users, particularly those who identify themselves as strong political partisans. If exposure to social media posts from Russian foreign influence accounts during the 2016 US election was similarly concentrated, their impact on changing candidate preferences may have been minimal. In the attempt to verify this hypothesis, Eady et al. [23] combined US longitudinal survey data from over 1496 respondents with Twitter data. They found that exposure to the Russian foreign influence campaign was heavily concentrated among a small fraction of users who identified themselves as Republicans. Moreover, they found no evidence of a significant relationship between exposure to the campaign and changes in attitudes, polarization, or voting behavior in the 2016 US election.
While prior research extensively explored the influence of IRA accounts on individuals’ voting intentions and Twitter discussions, it provided limited attention to the broader set of suspended accounts not flagged as IRA. During the ”Twitter purge” in May 2018, Twitter suspended numerous accounts, including those unrelated to the IRA [24, 25], with IRA accounts constituting only a smaller fraction of this overall set.
Considering that all these accounts faced suspension for violating Twitter rules, it prompts the question of whether, beyond IRA accounts, other accounts might have attempted to influence Twitter discourse and more broadly the 2016 US election.
Our findings provide evidence supporting this notion. A consistent group of suspended accounts exhibits similarity with IRA accounts in terms of the information they interact with and disseminate to the broader Twitter community. We demonstrate that the group of suspended accounts did indeed influence, in a Granger-causal manner, the retweet activity of undecided users and weak supporters—individuals uncertain about their voting decisions—in terms of political polarization.
The paper is organized as follows: Section 2 outlines the data collection and analysis methods utilized throughout the study. In Section 3, we present the findings of our research, which include: a) the characterization of users within our dataset based on the content they share; b) the use of the IRA ego network as a means to identify a group of suspended users with similar behavioral patterns; and c) an evaluation of how this specific group of suspended accounts influences Twitter discourse, utilizing Granger causality for assessment. The manuscript concludes with a thorough discussion and conclusion section, summarizing the key insights gained from our analysis.
2 Methods
2.1 Dataset
In this study, we combine the IRA dataset with a dataset containing tweets posted between June 1st and election day, November 8th, 2016. The data were collected continuously using the Twitter search API with the names of the two presidential candidates [3, 26, 27]. The 2016 dataset consists of 171 million tweets sent by 11 million users.
On the other hand, from June 1st to November 8th, 2016, 556 IRA accounts published 391680 tweets in English. According to [17], the content of these tweets aimed to sow and amplify political and social discord in the United States and manipulate the 2016 American presidential election. See Supplementary Sections 1 and 2 for more information.
To retrieve the account status of each user in the 2016 dataset, we used the Twitter users API, as of October 2023. It allows us to classify each account as suspended, not found, not verified, or verified. On Twitter, a suspended account refers to an account that has been temporarily or permanently disabled by Twitter due to a violation of its rules or policies. In contrast, a not found account is not deleted by Twitter but is no longer available because the user has chosen to delete or deactivate it. A not verified account on Twitter is an account that has not been officially confirmed by Twitter. Verification is a process through which Twitter verifies the authenticity and identity of notable public figures, organizations, or brands. On the other hand, a verified account on Twitter has undergone the verification process and has been confirmed by Twitter as an authentic representation of a notable public figure, organization, or brand. Verified accounts are distinguished by a blue checkmark badge next to their username, indicating their credibility and authenticity. Important to note that this is no longer the case (as of November 29th, 2023), as now anyone can buy the blue checkmark.
Among the 11 million users, 73.8% are not verified, 17.7% are not found, 7.7% are suspended, and 0.8% are verified. In this dataset, the IRA accounts account for less than 1% of the users (554 accounts).
2.2 News categories
In order to control for the type of information under analysis, we focus on tweets that contain at least one URL (Uniform Resource Locator) pointing to a news website outside of Twitter. We classified URL links for outlets that mostly conform to professional standards of fact-based journalism in five news media categories: right, right leaning, center, left leaning, and left. The classifications rely on the website allsides.com (AS), followed by the bias classification from the website mediabiasfactcheck.com (MBFC) for outlets not listed in AS (both accessed on 7 January 2021 for the 2020 classification) [3, 26, 27, 22]. We also include three additional news media categories to include outlets that tend to disseminate disinformation: Extreme bias right, Extreme bias left, and Fake news [3, 26, 27, 22]. Websites in the fake news category have been flagged by fact-checking organizations as spreading fabricated news or conspiracy theories, while websites in the extremely biased categories have been flagged for reporting controversial information that distorts facts and may rely on propaganda, decontextualized information or opinions misrepresented as facts. Supplementary Table 1 offers the list of news outlets per category considered in this work.
In the 2016 dataset, 2.3 million users shared 30.7 million tweets that contained URLs directing to news outlets. In the IRA dataset, 334 IRA accounts posted 23,806 tweets that included hyperlinks to news outlets.
2.3 Retweet network
In the context of a news category network, a link between two users occurs every time a user $u$ retweets a user’s tweet $v$ that contains a URL linking to a website belonging to one of the news media categories. The direction of the connection goes from $v$ to $u$, i.e., the direction of the information flow between Twitter users. We do not include multiple links in the same direction between the same two users, nor do we include self-links. The degree of a node in the network is defined as the number of edges connected to it. The out-degree of a node $u$, $k_{out}^{u}$, represents the number of unique users who retweeted $u$. On the other hand, the in-degree of a node $u$, $k_{in}^{u}$, represents the number of users retweeted by node $u$. It is worth noticing that, by construction, these networks are balanced directed networks, and as such, $\langle k_{in}\rangle=\langle k_{out}\rangle=\langle k\rangle/2$.
When building the ego networks, we proceed similarly. We construct separate networks for each of the four types of interactions: retweeting, mentioning, replying, and quoting. Each node in the network represents either an IRA or a non-IRA user. A link between two users occurs every time a user $u$ interacts with a user $v$ through a type of interaction. The direction of the connection goes from $v$ to $u$, i.e., the direction of the information flow. Connections are allowed between IRA nodes and between IRA and non-IRA nodes. We do not consider interactions among non-IRA nodes. We consider multiple interactions between two users; that is, networks are weighted by the number of times users interact.
Starting from the four interaction networks, we build an aggregated network (referred to as IRA ego network or IRA aggregated ego network, interchangeably) by considering all types of interaction and removing self-loops. As usual, an edge connecting node $u$ with node $v$ means there was at least one type of interaction between them. The edge is weighted by the number of interactions among $u$ and $v$. The directions of the links are according to the flow of information. The resulting structure is a directed weighted network of 179,783 nodes (524 of which are IRA accounts) and 432,429 edges (see Table 2).
2.4 Sampling strategies
To avoid sample bias, we randomly extracted the same amount as the number of IRA users for each group and category (making sure not to select the IRA users). We average the in/out degree over 1000 realizations. For each realization, sampling was without replacement. We refer to them as $(\overrightarrow{k_{type}^{s}}^{i},\overrightarrow{k_{type}^{s}}^{j})$, where the superscript $s$ indicates that the degree considered comes from the sampled nodes. Supplementary Table 7 displays the sampled average degree for each group in each news category, together with the standard error. Refer to Supplementary Table 8 for a view of the non-sampled case.
2.5 Two-sample Kolmogorov-Smirnov test
To test for differences in the in/out-degree activity of suspended, not found, not verified, verified, and IRA accounts, we employed a Two-sample Kolmogorov-Smirnov test with null hypothesis $H_{0}$: $F_{i}(x)=F_{j}(x)$ where $x=k_{type}^{s}$. The superscript $s$ indicates that the degree considered comes from the sampled nodes, $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). The null hypothesis, denoted as $H_{0}$, assumes that the activity of a given user in each interaction type, represented by $x=(x_{out},x_{in})$, follows the condition $F_{out}(x)=F_{in}(x)$ for every $x$. Here, $F_{out}(x)$ and $F_{in}(x)$ represent the cumulative density functions (CDF) for the “out” and “in” directions, respectively. The alternative hypothesis, on the other hand, suggests that $F_{out}(x)<F_{in}(x)$ (or $F_{out}(x)>F_{in}(x)$) for at least one $x$.
It is worth noting that these hypotheses describe the CDFs of the underlying distributions, not the observed data values. For example, suppose $x_{out}\sim F_{out}$ and $x_{in}\sim F_{in}$. If $F_{out}(x)>F_{in}(x)$ for all $x$, the values $x_{out}$ tend to be less than $x_{in}$. We set a level of 5%, meaning that we will reject the null hypothesis and favor the alternative if the p-value is less than 0.05.
2.6 Supporters identification
Similarly to [3], we use a supervised classifier to classify each tweet in favor of Donald Trump or Hillary Clinton. The training set was built using the hashtag co-occurrences network to investigate Twitter users’ opinions on the two presidential candidates. We classified a user as a supporter of Trump if the number of her/his tweets supporting Trump $N_{\textit{pro-T}}$ is greater than the number of tweets supporting Clinton $N_{\textit{pro-C}}$. We define the support of a given user toward the candidates as $S=N_{\textit{pro-T}}-N_{\textit{pro-C}}$. If $S>0$, the user supports Trump. Otherwise, the user is likely to support Hillary. The highest the value of $S$ in absolute terms, the strongest the support. Considering all the users in the dataset, 65% of them support Hillary Clinton while 28% are in favor of Donald Trump (7% are unclassified as they have the same number of tweets in each camp) [3]. When considering only the users interacting with IRA accounts, 25% of the users are classified as Clinton supporters, and 72.6% of the users are classified as Trump supporters.
2.7 Supporting classes
To distinguish between strong and weak supporters based on their $S$ values, we utilize the interquartile range (IQR) of $S$, defined as $IQR=Q_{3}-Q_{1}$, where $Q_{3}$ represents the third quartile and $Q_{1}$ represents the first quartile. In this analysis, a positive value of $S$ indicates a likelihood of supporting Trump, while a negative value of $S$ suggests a preference for Clinton. The magnitude of $S$ quantifies the degree of support for a particular candidate. Users who consistently retweet in favor of Trump referred to as strong supporters, exhibit higher $S$ values. Conversely, users with significantly negative $S$ values can be associated with strong supporters of Clinton. Users with $S=0$ are categorized as undecided since they display an equal number of tweets supporting both candidates.
In addition to the undecided category, we define four classes of supporters based on the interquartile range (IQR) of $S$ values. For Trump supporters, the IQR is calculated over the values of $S>0$, while for Clinton supporters, the IQR is computed using the absolute values of $S<0$. We identify weak Trump (Clinton) supporters as users whose $S$ values fall below $Q_{3}+1.5IQR$. On the other hand, strong Trump (Clinton) supporters are individuals whose $S$ values exceed $Q_{3}+1.5IQR$. This classification scheme allows us to distinguish between different levels of support.
Alternatively, we can consider the entire distribution of $S$ and define strong Trump supporters as users with $S$ values above $Q_{3}+1.5IQR$. Similarly, strong Clinton supporters are those with $S$ values below $Q_{1}-1.5IQR$. Weak Trump supporters fall within the range $Q_{1}-1.5IQR\leq S\leq Q_{3}+1.5IQR$ with $S>0$, while weak Clinton supporters fall within the same range but with $S<0$. This alternative classification, in the case of users interacting with IRA, results in 12.7% of users identified as strong Trump supporters, 60% as weak Trump supporters, 2.6% as strong Clinton supporters, and 22.5% as weak Clinton supporters. These percentages slightly differ from the ones obtained using the other approach mentioned in the main paper.
3 Results
3.1 Accounts characterization
Our analysis commences with a general characterization of the accounts active on the Twitter platform around the topic “election” during the 2016 US presidential elections. Refer to the Methods section, specifically Section 2.1 for a comprehensive description of the dataset. Users in this dataset are classified into distinct groups, encompassing IRA-flagged accounts by Twitter, along with not found, not verified, verified, and suspended accounts. This section aims to characterize the various account groups in terms of the information they spread on Twitter and draw comparisons with the IRA-flagged accounts. The rationale for using IRA as a benchmark in our analyses, as explained in the introduction, is to assess the impact of other accounts displaying behavior similar to IRA, specifically those favoring the right political candidate.
(a)Refer to caption
(b)Refer to caption
Figure 1: Distribution of tweets and clients’ type per account type. (a) The fraction of tweets with a URL pointing to a website belonging to one of the categories. Normalization is computed per group, meaning that, for example, the fraction of not verified tweets per category sums up to one. In each of the five groups, the order of bars is kept the same. We always display bars in the following orders: fake news, extreme bias right, right, right leaning, center, left-leaning, left, and extreme bias left. See Supplementary Table 2 for further details. (b) The fraction of tweets with a URL pointing to a website belonging to one of the categories posted through non official sources. Refer to Supplementary Table 3 for a complete view of the percentages.
The initial distinction among account types concerns the kind of information users engage with (see Section 2.2). Verified accounts, as shown in Fig. 1a, have a higher fraction of center and left-related (left and left-leaning) tweets, while IRA, suspended, and not found accounts exhibit a higher fraction of right-related tweets. Unlike IRA accounts, which show a significant percentage of center and left-leaning related news, suspended accounts have the lowest fraction of left-related content and the highest fraction of fake-related content. Supplementary Tables 2 and 3 provide a full breakdown of these percentages.
In Fig. 1b, we show the percentage of tweets shared through non official clients for each media category. To ensure comparability, we normalize the percentages per account type by the total group activity, including both official and non official clients. For official client details, refer to Supplementary Table 5. Analyzing tweet clients offers valuable insights into tweet origins, especially their potential bot-generated nature. Non official clients, encompassing applications like ifttt and dlvrit, span professional automation tools to manually programmed bots.
Figure 1b shows that verified accounts exhibit the highest fraction of tweets from non official clients, constituting 22.9% of their total activity. The most frequently used clients for verified accounts, such as Hootsuite and Socialflow, are renowned for automating interactions within the Twitter ecosystem. These verified accounts, often belonging to journalists or public figures, utilize such tools for social media activities. Suspended accounts rank second at 23.9%, with Dlvrit as their primary client, closely followed by IRA accounts at 22.9%, mostly relying on Twitterfeed for automated activity. Not verified and not found accounts exhibit below 13% non-official client usage, with Twitterfeed being the most used client. It is noteworthy that verified accounts predominantly used non official clients to disseminate center, left-leaning, and left news. In contrast, IRA accounts utilized non official clients mainly for right-related content, with a smaller percentage of left-leaning material. Suspended accounts employed non-official clients primarily for fake news dissemination.
Figure 2: Share of original tweets per group. The percentage of original tweets per group and news category is presented, with normalization conducted over the total activity of each group. This means that the sum of the percentages per group represents the total fraction of original tweets. Further details can be found in Supplementary Table 6.
Figure 2 shows the proportion of original tweets shared by each group and category. The normalization is over the total activity of each account type, meaning that the sum of the percentages per each account type represents the total fraction of original tweets. The group with the highest share of original tweets is the verified one, with a value of 71.2%. This group also shows the lowest share of original tweets linking to fake news and the highest share of original tweets related to the center, left-leaning, and left categories. IRA accounts instead show together with not found accounts, the lowest share of original tweets, with a percentage of 29.4% and 27.3%, respectively. Most of the original tweets shared by IRA belong either to the left leaning or center categories. Not found accounts have a more homogeneous distribution of original tweets among the different categories. Suspended accounts, with 38.6% of original tweets, show the highest percentage of original tweets related to the fake category and the extreme bias right category. Not verified accounts (32% of original tweets) show higher percentages in the center and left leaning.
Finally, we test whether suspended, not found, verified, not verified, and IRA behave differently in terms of their in/out activity in each news category network (see Section 2.3 and Table 1). We employ a two-sample Kolmogorov-Smirnov test [28, 29] (two-sided version, see Section 2.5) with null hypothesis $H_{0}$: the data are drawn from the same distribution. We performed the test for each two-pair combination of the groups $(\overrightarrow{k_{type}}^{i},\overrightarrow{k_{type}}^{j})$, with $i,j\in$ (suspended, not found, verified, not verified, IRA), and $type\in$ (in, out). The $\overrightarrow{k_{type}}$ vector contains the values of $\langle k_{type}\rangle$ for each category network. We adapt sampling strategies to avoid sampling bias error, as explained in Section 2.4. Refer to Supplementary Table 7 for a view of the sample degrees. While the news category networks considered here are not weighted, we did check for differences in the weighted case and could not find any significant variations.
Figure 3: Two-sample Kolmogorov-Smirnov tests. Results of the two sample Kolmogorov Smirnov test between groups based on their out degree. Yellow boxes indicate rejection of the null hypothesis in favor of the two-sided alternative, suggesting that the data were not drawn from the same distribution. We observe that $H_{0}$ is only rejected between IRA and suspended accounts, indicating a similarity in the out-degree activity of these two groups. Furthermore, verified accounts exhibit an out-degree activity that differs from that of the other groups.
Figure 3 shows the results of the tests for the out-degree and in-degree, respectively. We used a heatmap representation, where the yellow color indicates the rejection of the null hypothesis $H_{0}$ in favor of the default two-sided alternative, suggesting that the data were not drawn from the same distribution.
In comparing out-degree activity, verified accounts consistently exhibit distinct behavior from other groups, rejecting the null hypothesis. IRA accounts, however, display similar behavior to suspended accounts while differing from verified, not found, and not verified accounts. In terms of in-degree results, verified accounts differ from suspended, not found, and not verified accounts, but align with IRA users.
Among the various groups of accounts, our analyses reveal that IRA accounts and suspended accounts share similar interests in terms of the news outlets they reference, with suspended accounts showing a higher interest in fake and extreme bias right related content. Both groups also demonstrate a similar use of non-official clients, though with differences in the information transmitted through them. Not very dissimilar are not found accounts, which, however, display very low usage of non official clients.
Notably, our analyses uncover parallels in the behavior of IRA and suspended accounts concerning out-degree activity in each category. These resemblances might signify a shared effort by suspended accounts to steer Twitter discourse toward the right political agenda. However, it is crucial to emphasize that while this similarity with IRA does not imply collaboration, it is highly improbable that the entire set of suspended accounts is involved in this endeavor. Therefore, identifying a representative subset of suspended accounts participating in this intent becomes imperative.
3.2 IRA ego network
To pinpoint a representative subset of suspended accounts employing strategies akin to IRA accounts, we look into the IRA ego network. This network encompasses all users interacting with IRA through retweets, mentions, quotes, or replies (see Section 2.3). Table 2 summarizes the key features of these networks.
We analyze the interactions of users with different account types, such as verified, not verified, suspended, and not found. Specifically, we focus on the top 1000 users involved in each interaction type (retweet, reply, quote, and tweet) by considering both in and out degrees. For instance, in the case of retweet interaction, we consider the top 1000 (most retweeted) non-IRA users who were retweeted by IRA accounts in the “out” direction. Similarly, in the “in” direction, we examine the top 1000 (most retweeting) non-IRA users who retweeted IRA accounts.
(a)Refer to caption
(b)Refer to caption
Figure 4: Account status: top 1000 active users. We show the fraction of the top 1000 active users per status (in %) in each interaction network (retweet in blue, mention in red, reply in green, and quote in gray). Possible statuses are: Verified, Suspended, Not Verified and Not Found. (a) “out” direction: users are retweeted, mentioned, quoted, or replied to by IRA users. On average, 18.4% of these users have verified accounts, 49.4% are not verified, 18.9% are suspended, and 11.4% are not found accounts (b) “in” direction: users retweet, mention, quote, or reply to IRA users. On average, 45.1% of them are not verified, 35.7% is suspended, and 17.1% is not found.
Figure 4a displays the distribution of the users interacted by IRA (connections go from the non-IRA users to IRA) into the different account types. On average, 18.4% of these users have verified accounts, 49.4% are not verified, 18.9% are suspended accounts, and 11.4% are not found accounts. Notably, among the verified accounts, we identified the official profile of President Donald Trump and popular news outlets such as The Guardian and FOX NEWS. See Supplementary Tables 9, 10, 11, and 12 for a list of the top 20 accounts. Figure 4b, displays the distribution of the users who interact with IRA (connections go from the IRA to non-IRA users). None of the top 1000 users who engage with IRA have verified accounts. On average, 45.1% of them are not verified, 35.7% is suspended, and 17.1% is not found. See Supplementary Tables 13,14, 15 and 16 for a list of the top 20 accounts.
It is worth noting that the number of suspended accounts interacting with IRA (amounting to 30,622) is nearly 60 times larger than the number of IRA accounts.
It is noteworthy to observe that the majority of active users interacting with IRA are suspended users. Moreover, given that the most common type of interaction is retweeting, as indicated in Table 2, this suggests that most users tend to retweet IRA, likely with the intent to disseminate similar types of information. This ultimately confirms the notion that suspended accounts and IRA share similar views.
Ego polarization Analyzing user preferences in terms of political orientations (see Sections 2.6 and 2.7) reveals, as expected, a higher presence of right supporters in the IRA ego network. Specifically, 8.1% of users are strong supporters of Trump, 4% are strong supporters of Clinton, 64.5% are weak supporters of Trump, 21.1% are weak supporters of Clinton, and the remaining 2.3% of users are categorized as undecided. It’s important to note that when calculating user polarization in computing, we consider all users, regardless of their account status. The majority of suspended accounts are classified as Trump supporters, either weak or strong. This is detailed in Table 3, where we also provide the percentages broken down by account status.
This aligns with the notion that this subset of suspended accounts was oriented toward the right agenda. Additional details about the users’ classifications can be found in Supplementary Section 4.
It is also worth noticing that we found a substantial number of IRA accounts classified as weak Clinton supporters. This suggests a dual strategy employed by the IRA in their campaign, as already suggested in the existing literature [18, 30]. One aspect of this strategy involves reinforcing the opinions of users classified as strong Trump supporters. On the other hand, another set of IRA accounts aims to expand their reach within left-leaning accounts by mentioning verified accounts classified as weak and strong Trump supporters.
Community structure The exploration of this dual strategy and the role played by suspended accounts in it continues through the identification of communities within the aggregated IRA ego network. This network is formed by merging the networks for the four types of interaction, as detailed in Section 2.3.
We perform multiscale community detection to the largest connected component of the undirected weighted version of the aggregated network, which contains 99.9% of the initial nodes. To assess community stability, we utilize Markov stability, as detailed in Supplementary Section 5.
We reveal the existence of two prominent communities, a finding that adds intrigue as it aligns with the notion that the IRA orchestrated two distinct micro-campaigns. These communities reflect the polarization of users. In the biggest community, 76.6% of users are classified as weak Trump supporters and 9.5% as strong Trump supporters, while the second biggest community left has 70.2% of weak Clinton supporters users and 13% of strong Clinton supporters users, see Table 4. These results align with the existing literature [18, 30]. Moving forward, we will refer to these communities as the right and left communities.
The distinction in the political orientation of these two communities is further supported by analyses of the hashtag clouds constructed from the content shared by users, as discussed in Supplementary Section 6. Users in the right community tend to share hashtags in support of Trump and against Clinton, while the opposite holds true for the left community.
Within the right community, suspended accounts comprise 20% of the total nodes in the community. This percentage decreases by more than half in the left community (refer to Table 4). This outcome suggests a difference in strategies between suspended accounts and IRA, with IRA implementing a dual strategy (targeting both right and left users), while suspended accounts predominantly focus on targeting right-oriented users. These suspended users account for 21.7% of the interactions, with the directions of the connections going from IRA to suspended accounts, as indicated in Supplementary Figure 7 and Supplementary Table 20. In the case of the left community, the type of interaction is inverted, and in most cases, it is the IRA engaging with the other groups, particularly not verified and verified, with the connections going from suspended to IRA accounting for 12.5% of the overall connections in the community.
The above-presented results demonstrate that utilizing the IRA ego network as a proxy to identify a contained group of suspended accounts aligning with the right ideologies is an effective strategy. We find that the majority of these suspended users are classified as Trump supporters. By examining the dual strategies employed by the IRA campaign, we reveal that the majority of suspended accounts were following the strategies of the right community, specifically targeting Trump supporters. The directions of the connections suggest that this was primarily done through interactions with IRA accounts, including retweets, mentions, quotes, and replies.
Expanded ego network The IRA ego network serves as a proxy for identifying suspended accounts that share similarities with IRA accounts. However, this does not ensure that this subset covers all suspended accounts involved in promoting right-related content. Additional suspended accounts may exist in the dataset, but rather than interacting directly with IRA, they might be interacting with the suspended accounts that are engaging with IRA. To investigate this possibility, we expanded the IRA aggregated network from the earlier section by incorporating interactions involving this subset of suspended nodes and all other users. In essence, we created the “suspended+IRA” ego network. This network, akin to the previous section, was constructed based on the four types of Twitter interactions. Supplementary Table 17 provides comprehensive information regarding each interaction network.
The inclusion of both IRA accounts and suspended accounts significantly amplifies the dimension of the aggregated network, increasing the number of nodes from 179,783 in the IRA ego network to 1,723,477 in the expanded ego network. This expanded network exhibits 45 times more connections than the aggregated ego network, with an average degree of $\langle k\rangle=11$. Similar to the IRA ego network, retweeting and mentioning interactions remain the most common types of interactions in the expanded network.
Next, we performed a multi-scale community detection analysis and explored different parameter values to identify the optimal partition. The resulting partition (resulting in two communities) preserves the communities’ polarization, as shown in Table 5, with the two expanded communities being mostly composed of supporters of Trump and Clinton. When scrutinizing the composition of these two communities, it becomes evident that they are predominantly comprised of not verified, verified, and not found accounts. Suspended accounts, despite constituting a smaller percentage compared to the IRA ego network, also exhibit significantly lower activity. Their number of connections in both directions represent less than 5% of the total connections, as illustrated in Table 5.
These findings indicate minimal and negligible interactions between suspended accounts within the IRA ego network and other suspended accounts. This emphasizes that the suspended accounts identified in the IRA ego network represent the most significant group of suspended accounts disseminating right-related information, similar to IRA, in the Twitter discourse.
3.3 Causal network patterns: IRA nodes versus suspended nodes
This section investigates the impact that suspended accounts and IRA have on shaping the Twitter discourse during the 2016 US presidential elections. Specifically, we scrutinize the causal relationships between IRA (and suspended accounts) tweet activity and the activity of supporting classes, namely weak Trump supporters, weak Clinton supporters, strong Trump supporters, strong Clinton supporters, and undecided users.
We employ a multivariate Granger causal network reconstruction approach to establish links between the activity of IRA (suspended) nodes and the supporting classes. This is achieved using the causal discovery algorithm [31, 32, 33], which tests the independence of each pair of time series for several time lags conditioned on potential causal parents using a Partial Correlation Independence test and it removes spurious correlations. We use the algorithm for variable selection and perform a linear regression using only the true causal link discovered. We choose linear causal effects for their reliability and interpretability, which allows us to compare causal effects as first-order approximations, estimate the uncertainties of the model, and construct a causal-directed weighted network [34]. The causal effect between a time series $X^{i}$ and $X^{j}$ at a time delay $\tau$, $I^{CE}_{i\rightarrow j}(\tau)$, is determined by the expected value of $X^{j}_{t}$ (in units of standard deviation) if $x^{i}(t-\tau)$ is perturbed by one standard deviation [34, 26]. However, an assumption of causal discovery is causal sufficiency, which assumes that every common cause of any two or more variables is present in the system [31]. In our case, causal sufficiency is not satisfied because Twitter’s activity is only a part of a larger social system. Therefore, the term “causal” should be understood as relative to the system under study [26].
We created time series of Twitter activities by counting the number of tweets posted by each node belonging to one of the supporting classes at a 15-minute resolution. We only consider users that belong to the verified and not verified classes, and only consider the tweets coming from official clients. Instead, for the IRA (suspended) nodes, we consider all the tweets, no matter the clients. To remove trend and circadian cycles from the time series, we utilized the STL (seasonal trend decomposition procedure based on Loess) method [35], which decomposes a time series into seasonal (in this case, daily), trend, and remainder components. We used the residuals of the STL filtering of the 15-minute tweet volume time series.
In simpler terms, Granger causality examines whether past retweet behaviors in one group can assist in forecasting the retweet behaviors of the other group. It doesn’t imply a direct cause-and-effect relationship but rather investigates whether changes in one group’s activity precede changes in the other group’s activity.
Tables 6 and 7 present the causal relationships among different groups in the two scenarios: one with only IRA nodes and the other with suspended nodes. The direction of each link is from the column group to the row group. For example, considering the strong Trump supporters, their causal effect on the weak Clinton supporters is measured at 0.16 $\pm$ 0.011, as shown in Table 6. The blue entries in the tables represent the auto-correlation of each time series. In both scenarios, the auto-correlations exhibit the strongest causal effects for all time series, except for the undecided group.
To identify the most significant causal links, a threshold of 0.16 (0.20 for the suspended scenario) was set on the causal relation, selecting connections that account for 75% of the total effect. These selected links are highlighted in bold in the tables. Figures 5a and 5b visualize the causal networks constructed using these connections. The nodes are colored as follows: dark red for strong Trump supporters, dark blue for strong Clinton supporters, light red for weak Trump supporters, light blue for weak Clinton supporters, orange for the IRA nodes, and gray for the undecided group. Arrows indicate the direction of maximal causal effect ($\geq$0.16 and $\geq$ 0.20) between two activity time series. The width of each arrow represents the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series.
(a) Refer to caption
(b) Refer to caption
Figure 5: Causal Networks. (a) Graph showing the maximal causal effects between the activity of the IRA nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.16$) between two activity time series. The width of each arrow is proportional to the strength of the causation, and the size of each node is proportional to the auto-correlation of each time series. Dark blue and dark red highlight the contribution of strong Clinton and Trump supporters, respectively. Light blue and light red are associated with the weak Clinton and Trump supporters, gray with the undecided users, and orange with the IRA nodes. The causal relation primarily flows from strong supporters of both Trump and Clinton to weak and strong supporters of opposing political candidates. Additionally, weak supporters from both sides play a role in influencing the undecided group, with weak Trump supporters receiving support from strong Trump supporters in their efforts. Notably, IRA nodes do not play a significant role in this causal network, suggesting that they have limited influence on shaping Twitter discourse. (b) Graph showing the maximal causal effects between the activity of the suspended nodes and the supporting classes of the presidential candidates. Arrows indicate the direction of the maximal causal effect ($\geq 0.2$) between two activity time series. Strong Trump supporters have a causal effect on Suspended nodes, which, in turn, have a causal influence on both weak supporters and the undecided group. Additionally, weak supporters continue to exert a causal effect on the undecided group. Strong Trump supporters have a causal effect on strong Clinton supporters, but not vice-versa.
Figures 5a and b, illustrate contrasting scenarios in terms of the causal network structure when considering IRA nodes alone versus suspended nodes. In Figure 5a, which represents the causal network considering IRA nodes only, the influence primarily flows from strong supporters of both Trump and Clinton to weak and strong supporters of opposing political candidates. Additionally, weak supporters from both sides play a role in influencing the undecided group, with weak Trump supporters receiving support from strong Trump supporters in their efforts. Notably, IRA nodes do not play a significant role in this causal network, suggesting that they have limited causation on users’ activity.
On the other hand, in Fig. 5b, which represents the causal network for suspended nodes, the structure shows substantial differences. Suspended nodes take on a central role, acting as a bridge between strong Trump supporters and the weak and undecided supporters. Strong Trump supporters have a causal effect on suspended nodes, which, in turn, have a causal influence on both weak supporters and the undecided group. Additionally, weak supporters continue to exert a causal effect on the undecided group. Interestingly, strong Trump supporters have a causal effect on strong Clinton supporters, but not vice-versa.
4 Discussions and Conclusions
Current research focuses on the role and impact of the Internet Research Agency (IRA) in the 2016 US presidential elections. This emphasis on IRA’s political interference may have overshadowed other campaigns with similar aims that were not linked to Russian origins. By merging the IRA public dataset with a collection of tweets spanning the five months leading up to the 2016 presidential elections, our objective is to investigate the presence and impact of suspended accounts—those not flagged as IRA—which might have contributed to the dissemination of content aligned with the Trump political agenda.
Our analysis reveals that the IRA and suspended accounts (not flagged as IRA by Twitter) do share many similarities, in terms of the type of news they share, the clients they use and the way participate in the Twitter social discourse, as highlighted in Section 3.1. However, expecting all the suspended accounts in our extensive dataset, comprising over 700,000 users (7.7% of the 10 million users, as detailed in Section 2.1), to exhibit such similarity is improbable. To pinpoint a representative group more akin to IRA, we leverage the IRA ego network.
Within the IRA ego network, we identified 30,622 suspended accounts, a number 60 times larger than the IRA accounts. These suspended accounts engaged through various interactions like retweeting, mentioning, replying, and quoting, with retweets and mentions being the most common. Aligning with existing literature [36, 37] that asserts the IRA aimed to support Donald Trump and sow discord in the U.S., we found that the majority of nodes in the ego network, including suspended accounts, are classified as Trump supporters (discussed in Section 3.2).
In the aggregated IRA ego network, approximately 2% of total users were directly exposed to IRA content, consistent with [23]. A multiscale community detection on this network revealed two communities, encompassing almost 90% of the total nodes, indicating user polarization. The larger community (community right) aimed to support Trump, while the smaller one (community left) interacted directly with Clinton supporters, potentially attempting to influence their opinions. This dual strategy aligns with current literature suggesting a multifaceted approach by the IRA [18, 30].
Contrary to the multifaceted strategy observed with the IRA, suspended accounts exhibit a more focused role in the right community. They contribute significantly, comprising over 20% of the connections in the community, with a majority of these connections originating from IRA to suspended accounts. Building on earlier findings, this suggests that the suspended accounts in this group were primarily engaged in mentioning and retweeting IRA content, likely aiming to inundate the social platform with ideas consistent with the right political agenda.
Having identified a group of suspended accounts resembling IRA behaviors (partially) and gaining insights into their intent, we proceed to measure the impact of both suspended accounts and IRA on shaping the Twitter discourse during the 2016 US presidential elections. This is achieved through the application of Granger causality to the tweet activity produced by IRA and suspended accounts, and each supporting class (refer to Section 3.3).
Our causal analyses reveal that the group of IRA accounts did not have a significant impact in influencing the candidate’s supporters, as shown in Figure 5a and detailed in [23]. However, the situation becomes more intricate when we consider the suspended accounts. We find that these users wielded a substantial influence on individuals categorized as undecided or weak supporters, potentially with the intention of swaying their opinions. This effect is graphically portrayed in Figure 5b, illustrating the bridging effect that suspended nodes played between strong Trump supporters and the group of weak supporters and undecided individuals.
It’s important to note that while Granger causality suggests that past retweet behaviors in one group can aid in predicting the retweet behaviors of the other group, it doesn’t imply a direct cause-and-effect relationship. The determination of such causality goes beyond the scope of this study. Additionally, utilizing tweet activity provides insights into user behavior, but conclusions regarding changes in users’ vote intentions require longitudinal data. It is, however, noteworthy that the similarity observed between the Internet Research Agency (IRA) and the group of suspended users, coupled with the fact that suspended accounts influenced the activity of undecided users, opens up the possibility of a new scenario, such as potential cooperation between the IRA and the identified group of suspended users. It’s also conceivable that this group of suspended accounts was part of the IRA’s campaign and remained undetected by Twitter. However, this remains purely speculative, and further analysis and data are needed to draw more concrete conclusions.
Furthermore, the lack of detailed information about the nature of suspended accounts, such as whether they are trolls or bots, is a limitation. While all possibilities are considered, the logistical challenges of controlling a group of over 30,000 accounts make it more likely that this set of suspended accounts predominantly consists of bots.
In summary, this study suggests a scenario in which a significant group of suspended accounts, often overshadowed by the IRA narrative, played a crucial role during the 2016 US presidential elections. Further research is required to better understand their impact on political user preferences.
List of abbreviations
IRA: refer to the Internet Research Agency.
Declarations
Availability of data and materials
The Twitter data are provided according to its terms and are available at https://osf.io/g4hws/ and https://github.com/makselab/IRA-and-suspended-accounts.. Analytical codes are available in the same repositories.
Competing interests
The authors declare that they have no competing interests.
Funding
HAM was supported by NSF-HNDS Award 2214217. JSA gratefully acknowledges the Brazilian agencies FUNCAP, CNPq and CAPES, the National Institute of Science and Technology for Complex Systems in Brazil, and the PRONEX-FUNCAP/CNPq Award PR2-0101-00050.01.00/15 for financial support. MS gratefully acknowledges the Brazilian agency CAPES Award 88887.899221/2023-00. ZZ was supported by the National Natural Science Foundation of China under project No. 62302319 and R&D Program of Beijing Municipal Education Commission (Grant No. KM202210038002).
Authors’ contributions
HAM, AB and MS conceived the research; MS and AB designed and supervised the research; MS and AB coordinated and supervised the analysis; MS and ZZ performed the analyses. MS, ZZ, JSA, AB, and HAM analyzed the results; MS and ZZ wrote the first draft. All authors edited and approved the paper.
Acknowledgements
Not applicable.
Full Network IRA
N E $\langle k\rangle/2$ $N_{IRA}$ $\langle k_{out}\rangle$ $\langle k_{in}\rangle$
Fake 175,605 1,143,083 6.5 54 32.5 5.4
Extreme bias right 249,659 1,637,927 6.6 70 27.3 7.1
Right 345,644 1,797,023 5.2 84 30.1 4.7
Right leaning 216,026 495,307 2.3 67 6.3 1.7
Center 864,733 2,501,037 2.9 163 4.6 2.6
Left leaning 1,043,436 3,570,653 3.4 140 8.9 2.2
Left 536,903 1,801,658 3.4 105 4.3 2.4
Extreme bias Left 78,911 277,483 3.5 10 0.0 1.6
Table 1: Retweet categories’ networks. The table contains the characteristics of each of the eight retweet networks, such as the number of nodes $N$, the number of edges $E$, and the average degree $\langle k_{in}\rangle=\langle k_{out}\rangle=\langle k\rangle/2$. We also report the number of IRA users in each retweet network $N_{IRA}$, as well as their average in-degree $\langle k_{in}\rangle$ and out-degree $\langle k_{out}\rangle$.
Full networks IRA users
N E $\langle k\rangle$ $N_{IRA}$ $\langle k_{out}\rangle$ $\langle k_{in}\rangle$
Retweeting 154,366 360,265 2.3 497 468.4 262.0
Mentioning 70,926 197,644 2.8 508 353.1 41.7
Replying 14,225 16,775 1.2 193 71.2 16.4
Quoting 19,195 31,538 1.6 353 36.4 53.6
Aggregated 179,783 432,429 2.4 524 486.8 343.8
Table 2: Interactions ego networks. The table contains information about each type of interaction network, as well as information about their aggregated version. We report the number of nodes $N$, edges $E$, the average degree $<k>$, and the number of IRA nodes $N_{IRA}$, with their in/out-degree. Retweeting and mentioning are the two most frequent types of interactions between IRA and non-IRA users.
Strong Trump supporters Weak Trump supporters Strong Clinton supporters Weak Clinton supporters Undecided
Not Found 8.13 73.44 2.17 14.49 1.76
Not Verified 6.66 63.79 4.20 22.94 2.41
Suspended 14.43 67.93 3.56 11.78 2.30
Verified 0.25 15.84 11.63 68.95 3.33
IRA 0.32 34.62 0.00 60.26 4.81
Table 3: Groups classes. The table presents the distribution of different user types, namely not found, not verified, suspended, verified, and IRA users, among the various supporting classes. The percentages in the table are normalized per account type, meaning that the sum of percentages of a given account type for each supporting class adds up to 100%.
Community right Community left
n° nodes 135,846 24,318
n° IRA nodes 160 114
Undecided 1.8 % 4.1 %
Strong Trump supporters 9.5 % 2.1 %
Weak Trump supporters 76.6 % 10.5 %
Strong Clinton supporters 1.8 % 13.0 %
Weak Clinton supporters 10.3 % 70.2 %
Not Found 21.2 % 11.6 %
Not Verified 56.6 % 63.8 %
Suspended 20.3 % 9.6 %
Verified 1.7 % 14.5 %
IRA 0.1 % 0.5 %
Table 4: IRA ego network: partition characteristics. Characteristics of the communities in the IRA ego network. We display information for the communities with at least 10% of the nodes of the overall network. For each community, we report the number of nodes, the number of IRA accounts, the share of supporting classes and the distribution of users among different groups.
Community E-right Community E-left
n° nodes 718,825 314,559
n° IRA nodes 25,565 4,359
Undecided 5.4 % 4.1 %
Strong Trump supporters 10.2 % 2.6 %
Weak Trump supporters 55.9 % 16.0 %
Strong Clinton supporters 3.7 % 10.6 %
Weak Clinton supporters 24.8 % 66.6 %
Not Found 19.8 % 15.9 %
Not Verified 60.2 % 69.8 %
Suspended 14.2 % 8.9 %
Verified 2.2 % 4.0 %
IRA 3.6 % 1.4 %
Table 5: Expanded ego network: partition characteristics. Characteristics of the communities. We display information for the communities with at least 10% of the nodes of the overall network. For each community, we report the number of nodes, the number of IRA accounts, the share of supporting classes, and the distribution of users among different groups.
IRA Weak Trump supporters Weak Clinton supporters Strong Trump supporters Strong Clinton supporters Undecided
IRA 0.87 $\pm$ 0.006 0 0 0 0 0
Weak Trump supporters 0 0.54 $\pm$ 0.0155 0.08 $\pm$ 0.024 0.15 $\pm$ 0.015 0.16 $\pm$ 0.022 0.05 $\pm$ 0.019
Weak Clinton supporters 0 0.1 $\pm$ 0.021 0.78 $\pm$ 0.014 0.16 $\pm$ 0.011 0.13 $\pm$ 0.02 0.09 $\pm$ 0.02
Strong Trump supporters 0 0.01 $\pm$ 0.004 0.03 $\pm$ 0.015 0.47 $\pm$ 0.011 0.17 $\pm$ 0.015 0.01 $\pm$ 0.003
Strong Clinton supporters 0 0.06 $\pm$ 0.018 0.09 $\pm$ 0.013 0.16 $\pm$ 0.012 0.75 $\pm$ 0.013 0.07 $\pm$ 0.016
Undecided 0 0.26 $\pm$ 0.017 0.34 $\pm$ 0.021 0.16 $\pm$ 0.015 0.14 $\pm$ 0.02 0.2 $\pm$ 0.019
Table 6: Causal Links: IRA. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.
Suspended Weak Trump supporters Weak Clinton supporters Strong Trump supporters Strong Clinton supporters Undecided
Suspended 0.38 $\pm$ 0.012 0.05 $\pm$ 0.008 0.04 $\pm$ 0.012 0.27 $\pm$ 0.018 0.17 $\pm$ 0.020 0.03 $\pm$ 0.007
Weak Trump supporters 0.23 $\pm$ 0.027 0.52 $\pm$ 0.015 0.08 $\pm$ 0.021 0.10 $\pm$ 0.027 0.19 $\pm$ 0.023 0.05 $\pm$ 0.019
Weak Clinton supporters 0.29 $\pm$ 0.035 0.06 $\pm$ 0.017 0.79 $\pm$ 0.014 0.19 $\pm$ 0.037 0.12 $\pm$ 0.021 0.07 $\pm$ 0.019
Strong Trump supporters 0.05 $\pm$ 0.009 0.03 $\pm$ 0.011 0.02 $\pm$ 0.012 0.50 $\pm$ 0.011 0.15 $\pm$ 0.017 0.05 $\pm$ 0.012
Strong Clinton supporters 0.23 $\pm$ 0.030 0.02 $\pm$ 0.010 0.04 $\pm$ 0.012 0.20 $\pm$ 0.032 0.77 $\pm$ 0.012 0.04 $\pm$ 0.006
Undecided 0.24 $\pm$ 0.027 0.25 $\pm$ 0.015 0.34 $\pm$ 0.021 0.10 $\pm$ 0.027 0.16 $\pm$ 0.020 0.20 $\pm$ 0.019
Table 7: Causal Links: Suspended. We show the value of the maximal causal effect, $I_{i\to j}^{{\mathrm{CE,max}}}={\mathrm{max}}_{0<\tau\leq\tau_{{\mathrm{max}}}}\left|{I_{i\to j}^{{\mathrm{CE}}}(\tau)}\right|$ between each pair (i, j) of activity time series, where $\tau_{max}=18\times 15$min=4.5h is the maximal time lag considered, with standard errors. The arrows indicate the the direction of the causal effect. For each activity time series, we indicate in bold the three most important drivers of activity (excluding themselves). In blue, we highlight the auto-correlation of each node.